# AttentionStore: Cost-effective Attention Reuse across Multi-turn Conversations in Large Language Model Serving

### v2

@@ Abstract
Interacting with humans through multi-turn conversations is a fundamental feature of large language models (LLMs). However, existing LLM serving engines for executing multi-turn conversations are inefficient due to the need to repeatedly compute the key-value (KV) caches of historical tokens, incurring high serving costs. To address the problem, this paper proposes AttentionStore, a new attention mechanism that enables the reuse of KV caches (i.e., attention reuse) across multi-turn conversations, significantly reducing the repetitive computation overheads. AttentionStore maintains a hierarchical KV caching system that leverages cost-effective memory/storage mediums to save KV caches for all requests. To reduce KV cache access overheads from slow mediums, AttentionStore employs layer-wise pre-loading and asynchronous saving schemes to overlap the KV cache access with the GPU computation. To ensure that the KV caches to be accessed are placed in the fastest hierarchy, AttentionStore employs scheduler-aware fetching and eviction schemes to consciously place the KV caches in different layers based on the hints from the inference job scheduler. To avoid the invalidation of the saved KV caches incurred by context window overflow, AttentionStore enables the saved KV caches to remain valid via decoupling the positional encoding and effectively truncating the KV caches. Extensive experimental results demonstrate that AttentionStore significantly decreases the time to the first token (TTFT) by up to 87%, improves the prompt prefilling throughput by 7.8$\times$ for multi-turn conversations, and reduces the end-to-end inference cost by up to 70%. For long sequence inference, AttentionStore reduces the TTFT by up to 95% and improves the prompt prefilling throughput by 22$\times$.
@@ 1 Introduction
To reduce the serving cost and improve the inference performance, this paper proposes AttentionStore, a new attention mechanism that enables the reuse of KV caches (i.e., attention reuse) across multi-turn conversations rather than discarding them. When a conversation session becomes inactive, AttentionStore saves the corresponding KV cache in a KV caching system. Upon the resumption of the same session, AttentionStore loads and reuses the saved KV cache from the KV caching system, thereby eliminating the overhead of the repetitive computation. However, building such an efficient KV caching system for multi-turn conversations presents significant challenges.
@@
Firstly, the KV caching system serves as the external storage for GPUs and is attached to the GPUs via low-speed links. The use of the KV caching system brings about significant access overhead due to the need to transfer KV caches between HBMs and the KV caching system. The access overhead of KV caches is in the critical path of inference execution. This is because GPUs can only perform the computation of an inference job after successfully loading its corresponding KV cache into HBMs. Likewise, the subsequent inference jobs need to wait until the KV caches from the previous jobs are moved out of the HBMs if the HBM space is not enough. To reduce the KV cache loading overheads, AttentionStore uses a layer-wise pre-loading scheme to overlap the time of loading the KV cache with the inference computation layer by layer. To reduce the KV cache saving overheads, AttentionStore develops an asynchronous saving scheme that overlaps the time of saving KV caches with the inference computation.
@@
Secondly, the KV caches occupy a large amount of storage space that continuously expands during conversations. Prior works have attempted to reduce the inefficiency of repetitive KV computation by retaining the KV caches across multi-turn conversations in HBMs [19, 67]. However, this quickly exhausts the limited HBM capacity. We present an example of LLaMA-65B in Section 2.3, which shows the KV caches fully occupy the free space within the HBMs in 14 seconds. To address this challenge, AttentionStore explores and exploits slower but larger-capacity storage hierarchies than HBMs, including host memory and disks, to provide adequate storage space for caching KV caches.
@@
Thirdly, since disks have much larger capacity than the host memory (tens of TBs v.s. several hundreds of GBs), most KV caches are retained in disks for AttentionStore. As conversation requests arrive randomly, their corresponding KV caches are more likely to be located in disks, resulting in poor access performance. To address this problem, AttentionStore uses a scheduler-aware KV cache fetching scheme. This scheme pre-fetches the KV caches that are likely to be accessed from disks to the host memory, by utilizing the hints received from the interface job scheduler. When the free space of the host memory is not enough, AttentionStore also adopts a scheduler-aware eviction scheme to efficiently identify the most suitable KV caches in memory and evict them to disks or out of the system.
@@
Finally, when a conversation session surpasses the limit of the context window of LLMs, e.g., 4K in LLaMA-2 [49], LLMs generally truncate the oldest tokens and limit the context to the most recent tokens [33]. This truncation makes all saved KV caches of that conversation in AttentionStore invalid since the positional information of all tokens embedded in the KV cache is changed. To overcome this issue, AttentionStore decouples the positional encoding from the KV caches when saving them. It re-embeds the positional encoding into KV caches when loading them. After decoupling, truncation can be directly applied to the KV caches, thereby ensuring the reusability of the saved KV caches.
@@
We implement the AttentionStore and evaluate it using the real ShareGPT dataset [37]. Extensive experimental results demonstrate that AttentionStore significantly decreases the time to the first token (TTFT) by up to 87% and improves the prompt prefilling throughput by 7.8$\times$ for multi-turn conversations. It also reduces the end-to-end inference cost by up to 70%. For long sequence inference, AttentionStore reduces the TTFT by up to 95% and improves the prompt prefilling throughput by 22$\times$. To summarize, this paper makes the following contributions:
@@
  • We propose AttentionStore, a new attention that allows the reuse of the KV caches for any ensuing conversation turns of the same session, achieving a significant reduction in the recomputation overhead of KV caches in LLMs.
  • To improve the efficiency of AttentionStore, we design overlapped KV cache access, hierarchical KV cache placement, and positional encoding decoupled KV cache truncation schemes.
  • We thoroughly evaluate AttentionStore with real datasets to demonstrate its efficacy and efficiency.
@@ 2 Background and Motivation
This section begins with an overview of the fundamentals of generative LLM inference. It then delves into the inefficiencies that exist in LLMs during multi-turn conversations. The section ends with a discussion of the design opportunities for dealing with these inefficiencies and the challenges faced during the design of such a system.
@@ 2.1 Generative LLM Inference Basics
Transformer Architecture. The transformer has emerged as the widely accepted standard in generative LLM inference. The widely used LLMs like GPTs [32] and LLaMAs [48, 49] are built upon the autoregressive transformer architecture [51, 17]. During inference, these models process the prompt of the users and generate a response. The prompt is processed as a sequence of input tokens, and the response is generated by the model predicting the probability of subsequent tokens using the context of all the prior tokens. The transformer model consists of a chain of $l$ transformer layers. Each transformer layer is comprised of two steps, self-attention and feed-forward network (FFN).
@@
KV Cache: Within the entire process above, each token produces intermediate K and V tensors. When generating subsequent tokens, all KV tensors of preceding tokens are necessary for computing the self-attention. These K and V tensors are generally cached in GPUs, referred to as the KV cache. The KV cache typically has a large footprint. For example, GPT-3 [32, 11] generates a 4.5MB KV cache for each token. The size of KV cache linearly increases with the number of prompt tokens. A conversation session containing thousands of tokens will produce several GBs of KV cache.
@@ 2.2 Autoregressive Generation
The decoding phase. The decoding phase generates output tokens with autoregressive iterations. The decoding phase takes token $s+1$ and the KV cache $[1:s]$ from the prefilling phase as input to compute the KV cache $s+1$ and the token $s+2$. The generation process iteratively continues until the generated token is <eos> or the iteration number reaches the maximum allowed generation number. The decoding phase only happens sequentially due to the heavy data dependency on the previous iteration.
The two phases present significantly different characteristics in terms of execution time. The prefilling phase computes the KV cache in parallel. The duration of this phase is closely tied to the number of prompt tokens provided as input. As shown in Figure 1(b), the execution time of the prefilling phase increases as the number of input tokens grows. In contrast, the decoding phase only performs computation for a single token in each iteration, which makes the computation time for each iteration relatively constant.
@@ 2.3 Multi-turn Conversation Inference
Figure 3: Comparison of recomputation and AttentionStore.
@@
Figure 4: Recomputation inefficiencies. (a) The average numbers of historical tokens and new tokens in different turns of ShareGPT [37]. (b) The GPU time for prefilling all tokens and only new input tokens in ShareGPT with Mistral-7B [20] on 1 A100 GPU.
@@
However, executing multi-turn conversations in current LLM serving engines is inefficient due to the repetitive computation of KV caches across multiple conversation turns. As shown in Figure LABEL:fig:conversation, in the conversation turn 1, the LLM serving engine generates the KV cache of $q_{1}$ and $a_{1}$. After finishing turn 1, the LLM serving engine discards the KV cache to reclaim the HBM space. In turn 2, the LLM serving engine re-generates the KV cache of $q_{1}$ and $a_{1}$. In turn 3, the KV cache of $q_{1}$, $a_{1}$, $q_{2}$, and $a_{2}$ is re-generated. As the session expands, the historical tokens keep accumulating and the amount of repetitive computation significantly increases. As shown in Figure LABEL:fig:recompute-tokens, with the increase of the conversation turns, the percentage of historical tokens will be more than 99% in a new conversation. The repetitive computation time occupies 99% of the prefilling time (a.k.a., time to the first token) in the new conversation, as shown in Figure LABEL:fig:recompute-costs.
@@ 2.4 Opportunities and Challenges
Based on the analysis above, we observe that if the KV caches can be reused across multiple turns of conversations, up to 98% of prefilling cost can be reduced. Specifically, the KV caches of historical conversations can be saved in a KV caching system out of GPUs. Upon the reactivation of a conversation session, GPUs load the associated KV caches from the KV caching system and reuse them for the new-turn conversation. Nevertheless, to build an efficient KV caching system, there exist many significant challenges.
@@
2) High storage capacity requirement of KV caches. Storing the KV cache for each request consumes a substantial amount of storage space. For instance, when using 4 A100 GPUs each with 80GB HBM to run LLaMA-65B, prefilling 2K tokens consumes about 360 ms. This process generates 5GB of KV cache, indicating the generation speed of the KV cache is about 13.9GB/s. As 130GB of HBM space is allocated to store the model, the remaining 190GB of free HBM space will be fully occupied by the KV cache within 14 seconds. If spilling the KV cache to the host memory (e.g., 512GB space), the host memory will be filled in less than 1 minute. Using disks to save the KV cache can extend the storage space. However, this incurs worse access performance, as presented below.
@@
4) Unexpected invalidation of the saved KV caches. With the number of conversation turns increasing, the historical tokens can exceed the context window limitation. LLM serving engines generally perform token truncation [33, 16] to reduce the input prompt. The truncation has no impact on previous LLM serving engines since they always recompute the KV cache based on the input prompt following truncation. However, the truncation makes the KV caches saved in the KV caching system invalid, since the position of each token is changed after truncation. Thus it cannot match the old embedded positional encoding in the saved KV cache. Such context window overflow can occur with a high probability. As shown in Figure LABEL:fig:dist_tokennum, 47% and 30% of conversation sessions have a context longer than 2K and 4K, respectively. It means that when using the LLaMA-2 family with 4K context window [49], the context window overflow occurs in 30% of conversation sessions. When using the OPT family with 2K context window [63], the context window overflow occurs in 47% of conversation sessions.
3 The AttentionStore Design
@@ 3.1 Overview
We propose a new attention mechanism, called AttentionStore, to save the KV caches for all conversations, enabling the reuse of historical KV caches across multi-turn conversations, instead of discarding them as in conventional attention mechanisms. Specifically, AttentionStore saves the KV cache in a KV caching system when the associated conversation session is inactive. If the same conversation session is activated in the future, its KV cache is fetched from the KV caching system and reused for inference. By doing so, AttentionStore only executes partial prefilling, on just the new tokens input in the new turn of conversation, rather than prefilling all historical tokens. As shown in Figure LABEL:fig:cachedattention, when executing the inference of Turn 3, the KV cache of $q_{1}$, $a_{1}$, $q_{2}$, and $a_{2}$ is reused and only $q_{3}$ needs to be prefilled. AttentionStore effectively eliminates the repetitive computation overhead of the historical tokens, thereby reducing the prefilling cost.
Figure 5 shows the architectural overview of AttentionStore. It maintains a hierarchical KV caching system with efficient KV cache access, placement, and truncation techniques to address the challenges mentioned in Section 2.4.
Figure 5: The system architecture of AttentionStore.
For Challenge 1, to reduce the overhead of KV cache loading from the KV caching system into HBMs, AttentionStore leverages a layer-wise pre-loading scheme to overlap the KV cache loading with the inference computation. To reduce the KV cache saving overhead from HBMs to host memory, AttentionStore leverages an asynchronous saving scheme to overlap the saving with the inference computation. (§3.2).
For Challenges 2 and 3, to enlarge the available storage space for caching KV caches, AttentionStore employs multi-tier cost-effective storage mediums, i.e., host memory and disks. To reduce the impact of accessing slow disks on the inference performance, we present a scheduler-aware fetching scheme that leverages the hints from the job scheduler to prefetch KV caches to be accessed from disks to host memory. Meanwhile, to efficiently leverage the limited host memory space, we present a scheduler-aware eviction scheme that identifies the least valuable KV caches and evicts them to disks or out of the caching system (§3.3).
For Challenge 4, to deal with the invalidation of KV caches saved in AttentionStore due to context window overflow, we utilize a positional encoding decoupled truncation scheme to save the KV caches without positional encoding embedded, and hence support the truncation directly on KV caches. When loading the KV cache, AttentionStore re-embeds the new positional encoding into the KV caches (§3.4).

### v3

# Cost-Efficient Large Language Model Serving for Multi-turn Conversations with CachedAttention

@@ Abstract
Interacting with humans through multi-turn conversations is a fundamental feature of large language models (LLMs). However, existing LLM serving engines executing multi-turn conversations are inefficient due to the need to repeatedly compute the key-value (KV) caches of historical tokens, incurring high serving costs. To address the problem, this paper proposes CachedAttention, a new attention mechanism that enables reuse of KV caches across multi-turn conversations, significantly reducing the repetitive computation overheads. CachedAttention maintains a hierarchical KV caching system that leverages cost-effective memory/storage mediums to save KV caches for all requests. To reduce KV cache access overheads from slow mediums, CachedAttention employs layer-wise pre-loading and asynchronous saving schemes to overlap the KV cache access with the GPU computation. To ensure that the KV caches to be accessed are placed in the fastest hierarchy, CachedAttention employs scheduler-aware fetching and eviction schemes to consciously place the KV caches in different layers based on the hints from the inference job scheduler. To avoid the invalidation of the saved KV caches incurred by context window overflow, CachedAttention enables the saved KV caches to remain valid via decoupling the positional encoding and effectively truncating the KV caches. Extensive experimental results demonstrate that CachedAttention significantly decreases the time to the first token (TTFT) by up to 87%, improves the prompt prefilling throughput by up to 7.8$\times$ for multi-turn conversations, and reduces the end-to-end inference cost by up to 70%.
@@ 1 Introduction
To reduce the serving cost and improve the inference performance, this paper proposes CachedAttention, a new attention mechanism that enables the reuse of KV caches across multi-turn conversations rather than discarding them. When a conversation session becomes inactive, CachedAttention saves the corresponding KV cache in a KV caching system, named AttentionStore. Upon the resumption of the same session, CachedAttention loads and reuses the saved KV cache from AttentionStore, thereby eliminating the overhead of the repetitive computation. However, building such an efficient KV caching system for multi-turn conversations presents significant challenges.
@@
Firstly, AttentionStore serves as the external storage for GPUs and is attached to the GPUs via low-speed links. The use of AttentionStore brings about significant access overhead due to the need to transfer KV caches between HBMs and AttentionStore. The access overhead of KV caches is in the critical path of inference execution. This is because GPUs can only perform the computation of an inference job after successfully loading its corresponding KV cache into HBMs. Likewise, the subsequent inference jobs need to wait until the KV caches from the previous jobs are moved out of the HBMs if the HBM space is not enough. To reduce the KV cache loading overheads, CachedAttention uses a layer-wise pre-loading scheme to overlap the time of loading the KV cache with the inference computation layer by layer. To reduce the KV cache saving overheads, CachedAttention develops an asynchronous saving scheme that overlaps the time of saving KV caches with the inference computation.
@@
Secondly, the KV caches occupy a large amount of storage space that continuously expands during conversations. Prior works have attempted to reduce the inefficiency of repetitive KV computation by retaining the KV caches across multi-turn conversations in HBMs [18, 63]. However, this quickly exhausts the limited HBM capacity. We present an example of LLaMA-65B in Section 2.3, which shows the KV caches fully occupy the free space within the HBMs in 14 seconds. To address this challenge, CachedAttention explores and exploits slower but larger-capacity storage hierarchies than HBMs, including host memory and disks, to provide adequate storage space for caching KV caches.
@@
Thirdly, since disks have a much larger capacity than the host memory (tens of TBs v.s. several hundreds of GBs), most KV caches are retained in disks for CachedAttention. As conversation requests arrive randomly, their corresponding KV caches are more likely to be located in disks, resulting in poor access performance. To address this problem, CachedAttention uses a scheduler-aware KV cache fetching scheme. This scheme pre-fetches the KV caches that are likely to be accessed from disks to the host memory, by utilizing the hints received from the inference job scheduler. When the free space of the host memory is not enough, CachedAttention also adopts a scheduler-aware eviction scheme to efficiently identify the most suitable KV caches in memory and evict them to disks or out of the system.
@@
Finally, when a conversation session surpasses the limit of the context window of LLMs, e.g., 4K in LLaMA-2 [45], LLMs generally truncate the oldest tokens and limit the context to the most recent tokens [30]. This truncation makes all saved KV caches of that conversation in CachedAttention invalid since the positional information of all tokens embedded in the KV cache is changed. To overcome this issue, CachedAttention decouples the positional encoding from the KV caches when saving them. It re-embeds the positional encoding into KV caches when loading them. After decoupling, truncation can be directly applied to the KV caches, thereby ensuring the reusability of the saved KV caches.
@@
We implement the CachedAttention and evaluate it using the real ShareGPT dataset [33]. Extensive experimental results demonstrate that CachedAttention significantly decreases the time to the first token (TTFT) by up to 87% and improves the prompt prefilling throughput by up to 7.8$\times$ for multi-turn conversations. It also reduces the end-to-end inference cost by up to 70%. To summarize, this paper makes the following contributions:
@@
  • We propose CachedAttention, a new attention that allows the reuse of the KV caches for any ensuing conversation turns of the same session, achieving a significant reduction in the recomputation overhead of KV caches in LLMs.
  • To improve the efficiency of CachedAttention, we design overlapped KV cache access, hierarchical KV cache placement, and positional encoding decoupled KV cache truncation schemes.
  • We thoroughly evaluate CachedAttention with real datasets to demonstrate its efficacy and efficiency.
@@ 2 Background and Motivation
This section first introduces the fundamentals of generative LLM inference and then explores the inefficiencies in LLMs during multi-turn conversations. The section finally discusses the design opportunities for dealing with these inefficiencies and the challenges faced in designing such a system.
@@ 2.1 Generative LLM Inference Basics
Transformer Architecture. The transformer has emerged as the widely accepted standard in generative LLM inference. The widely used LLMs like GPTs [29] and LLaMAs [44, 45] are built upon the autoregressive transformer architecture [47, 16]. During inference, these models process the prompt of the users and generate a response. The prompt is processed as a sequence of input tokens, and the response is generated by the model predicting the probability of subsequent tokens using the context of all the prior tokens. The transformer model consists of a chain of $l$ transformer layers, each consisting of two phases: self-attention and feed-forward network (FFN).
@@
KV Cache: The above process produces intermediate K and V tensors for each token. When generating subsequent tokens, all KV tensors of preceding tokens are necessary for computing the self-attention. These K and V tensors are generally cached in GPUs, referred to as the KV cache. The KV cache typically has a large footprint. For example, GPT-3 [29, 10] generates a 4.5MB KV cache for each token. The size of KV cache linearly increases with the number of prompt tokens. A conversation session containing thousands of tokens will produce several GBs of KV cache.
@@ 2.2 Autoregressive Generation
[…]
The two phases present distinct execution time characteristics. The prefilling phase, computing the KV cache in parallel, has a duration closely tied to the number of input prompt tokens. As shown in Figure 1(b), the execution time of the prefilling phase increases with more input tokens. In contrast, the decoding phase computes for a single token per iteration, resulting in relatively constant computation time per iteration.
@@ 2.3 Multi-turn Conversation Inference
Figure 3: Comparison of recomputation and CachedAttention.
@@
Figure 4: Recomputation inefficiencies. (a) Average numbers of historical tokens and new tokens in different turns of ShareGPT [33]. (b) GPU time for prefilling all tokens and only new input tokens in ShareGPT with Mistral-7B [19] on 1 A100 GPU.
@@
However, executing multi-turn conversations in current LLM serving engines is inefficient due to the repetitive computation of KV caches across multiple conversation turns. As shown in Figure LABEL:fig:conversation, in conversation turn 1, the LLM serving engine generates the KV cache of $q_{1}$ and $a_{1}$. After finishing turn 1, the LLM serving engine discards the KV cache to reclaim the HBM space. In turn 2, the LLM serving engine re-generates the KV cache of $q_{1}$ and $a_{1}$. In turn 3, the KV cache of $q_{1}$, $a_{1}$, $q_{2}$, and $a_{2}$ is re-generated. As the session expands, the historical tokens keep accumulating and the amount of repetitive computation significantly increases. As shown in Figure LABEL:fig:recompute-tokens, with the increase of the conversation turns, the percentage of historical tokens will be more than 99% in a new conversation. The repetitive computation time occupies 99% of the prefilling time (a.k.a., time to the first token) in the new conversation, as shown in Figure LABEL:fig:recompute-costs.
@@ 2.4 Opportunities and Challenges
Based on the analysis above, we observe that if the KV caches can be reused across multiple turns of conversations, up to 99% of prefilling cost can be reduced. Specifically, the KV caches of historical conversations can be saved in a KV caching system out of GPUs. Upon the reactivation of a conversation session, GPUs load the associated KV caches from the KV caching system and reuse them for the new-turn conversation. Nevertheless, to build an efficient KV caching system, there exist many significant challenges.
@@
2) High storage capacity requirement of KV caches. Storing KV cache for each request consumes a substantial amount of storage space. For instance, when using 4 A100 GPUs each with 80GB HBM to run LLaMA-65B, prefilling 2K tokens consumes about 360 ms. This process generates 5 GB of KV cache, indicating the generation speed of the KV cache is about 13.9 GB/s. As 130 GB of HBM space is allocated to store the model, the remaining 190 GB of free HBM space will be fully occupied by the KV cache within 14 seconds. If spilling the KV cache to the host memory (e.g., 512 GB space), the host memory will be filled in less than 1 minute. Using disks to save the KV cache can extend the storage space but incur worse access performance, as presented below.
@@
4) Unexpected invalidation of the saved KV caches. With the number of conversation turns increasing, the historical tokens can exceed the context window limitation. LLM serving engines generally perform token truncation [30, 15] to reduce the input prompt. The truncation has no impact on previous LLM serving engines since they always recompute the KV cache from the truncated input prompt. However, the truncation makes the KV caches saved in AttentionStore invalid, since the position of each token is changed after truncation. Thus it cannot match the old embedded positional encoding in the saved KV cache. Such context window overflow can occur with a high probability. As shown in Figure LABEL:fig:dist_tokennum, 47% and 30% of conversation sessions have a context longer than 2K and 4K, respectively. It means that when using the LLaMA-2 family with 4K context window [45], the context window overflow occurs in 30% of conversation sessions. When using the OPT family with 2K context window [59], the context window overflow occurs in 47% of conversation sessions.
3 The CachedAttention Design
@@ 3.1 Overview
In this paper, we propose a new attention mechanism, called CachedAttention, which enables the reuse of KV caches across multi-turn conversations. Unlike the conventional attention mechanism that uses all prompt tokens for prefilling, CachedAttention uses the new tokens input in the new conversation turn and the KV caches of historical tokens for prefilling, as shown in Figure 3. Specifically, when the associated conversation session is inactive, CachedAttention saves the KV cache in a KV caching system, named AttentionStore, instead of discarding them as in the conventional attention mechanism. If the same conversation session is activated in the future, its KV cache is fetched from AttentionStore and reused for inference. By doing so, CachedAttention only executes the prefilling of partial prompt tokens, i.e., the new tokens input in the new turn of conversation, rather than prefilling all prompt tokens. As shown in Figure LABEL:fig:cachedattention, when executing the inference of Turn 3, the KV cache of $q_{1}$, $a_{1}$, $q_{2}$, and $a_{2}$ is reused and only $q_{3}$ needs to be prefilled. CachedAttention effectively eliminates the repetitive computation overhead of the historical tokens, thereby reducing the prefilling cost.
Figure 5 shows the architectural overview of CachedAttention. It maintains a hierarchical KV caching system, i.e., AttentionStore, with efficient KV cache access, placement, and truncation techniques to address the challenges mentioned in Section 2.4.
Figure 5: The system architecture of CachedAttention.
For Challenge 1, to reduce the overhead of KV cache loading from AttentionStore into HBMs, CachedAttention leverages a layer-wise pre-loading scheme to overlap the KV cache loading with the inference computation. To reduce the KV cache saving overhead from HBMs to host memory, CachedAttention leverages an asynchronous saving scheme to overlap the saving with the inference computation. (§3.2).
For Challenges 2 and 3, to enlarge the available storage space for caching KV caches, CachedAttention employs multi-tier cost-effective storage mediums in AttentionStore, i.e., host memory and disks. To reduce the impact of accessing slow disks on the inference performance, we present a scheduler-aware fetching scheme that leverages the hints from the job scheduler to prefetch KV caches to be accessed from disks to host memory. Meanwhile, to efficiently leverage the limited host memory space, we present a scheduler-aware eviction scheme that identifies the least valuable KV caches and evicts them to disks or out of the caching system (§3.3).
For Challenge 4, to deal with the invalidation of KV caches saved in CachedAttention due to context window overflow, we utilize a positional encoding decoupled truncation scheme to save the KV caches without positional encoding embedded, and hence support the truncation directly on KV caches. When loading the KV cache, CachedAttention re-embeds the new positional encoding into the KV caches (§3.4).